OpenStudy (anonymous):

Anybody know how to do this? At an all-you-can-eat barbecue fundraiser that you are sponsoring, adults pay 6$ for a dinner and children pay 4$ for a dinner. 212 people attend and raise 1128$. What is the total number of adults and the total number of children attending. (using systems of equations. a. what is a system of equations that you can solve this problem? b.what method would you use and why?

OpenStudy (anonymous):

Let x= adult price y=child's price

OpenStudy (anonymous):

How many adults and children? X+y=? Now do the prices ? x+4y= ?

OpenStudy (anonymous):

let: x= adult price y=child's price--->therefore x+y=212 and 6x+4y=1128---->therefore:x=212-y and y=212-x therefore since we know x and y, we substitute to get : 6(212-y)+4y=1128==>1272-2y=1128 therefore y=72

OpenStudy (anonymous):

same steps for x.

OpenStudy (anonymous):

or x+72=212 x=140